# Learning to Defer in Content Moderation: The Human-AI Interplay

### v2

@@ D.1 Comparison with Literature of Linear Contextual Bandits
Contrasting to linear contextual bandit, e.g. [APS11], and linear contextual bandit with delays, such as [VCP17, BXZ23], there are three main distinctions.
The first distinction is that those works use a different estimator $\hat{\boldsymbol{\theta}}$ by including all arrived data pointss into the dataset (unobserved labels are set as 00). In contract, our confidence set is constructed based on all reviewed posts in (13), which is a subset of all arrived posts endogenously determined by our algorithm.
The second distinction, as a result of the first, is that conditioned on the set of arrived posts, the matrix $\bar{V}_{t}$ in (13) is a fixed matrix in their settings, but is a random matrix in our setting because it is constructed from the set of reviewed posts. We thus must deal with intricacy of the randomness in the regression.
The third distinction lies in the analysis. The analysis of prior work crucially relies on an independence assumption, i.e., the event of observing feedback for a post within a particular delay period is independent from other posts, which enables a concentration bound on the number of observed feedback. In our setting, this is no longer the case as the delay in one post implies that other posts wait in the queue and thus delays could correlate with each other. We thus resort to the more intuitive estimator and analyze it via properties of our queueing systems.
@@ D.2 Regret guarantee of COLBACID (Theorem 4)
The proof uses [APS11, Theorem 2], which we restate in Fact 3. In particular, we recall that $Z_{t}=c_{\mathcal{S}(t)}=\boldsymbol{\phi}_{k(\mathcal{S}(t))}^{\intercal}\boldsymbol{\theta}^{\star}+\xi_{\mathcal{S}(t)}=\boldsymbol{X}_{t}\boldsymbol{\theta}^{\star}+\xi_{\mathcal{S}(t)}$. To apply Fact 3, it suffices to show that the sequence $\{\xi_{\mathcal{S}(t)}\}$ is conditionally $\sigma_{\max}-$sub-Gaussian for the filtration $\mathcal{F}_{t-1}=\sigma(\boldsymbol{X}_{1},\ldots,\boldsymbol{X}_{t},\xi_{\mathcal{S}(1)},\ldots,\xi_{\mathcal{S}(t-1)})$, i.e., $\mathbb{E}\left[e^{u\xi_{\mathcal{S}(t)}}\,\middle|\,\mathcal{F}_{t-1}\right]\leq\exp(u^{2}\sigma_{\max}^{2}/2)$ for any $u\in\mathbb{R}$.
Given that we assume the idiosyncratic noises $\{\xi_{j}\}$ of posts are independent samples, it holds that $\mathbb{E}\left[e^{u\xi_{\mathcal{S}(t)}}\,\middle|\,\mathcal{F}_{t-1}\right]=\mathbb{E}\left[e^{u\xi_{\mathcal{S}(\tau)}}\right]$. Since $\xi_{\mathcal{S}(t)}\sim\mathcal{N}(0,\sigma_{k}^{2})$ conditioned on the type of the post $k(\mathcal{S}(t))=k$, we have $\mathbb{E}\left[e^{u\xi_{\mathcal{S}(t)}}\right]=\exp(u^{2}\sigma_{k(\mathcal{S}(t))}^{2}/2)\leq\exp(u^{2}\sigma_{\max}^{2}/2)$, and thus $\{\xi_{\mathcal{S}(t)}\}$ is conditionally $\sigma_{\max}-$sub-Gaussian. Hence the conditions of Fact 3 are met and we finish the proof by applying Fact 3. ∎
@@
For $\delta\in(0,1)$, it holds that $\mathbb{E}\left[|\mathcal{T}_{E}|\right]\leq\frac{4B_{\delta}^{2}(T)\mathbb{E}\left[\sum_{t\in\mathcal{T}_{E}}\|\boldsymbol{\phi}_{k(j(t))}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}^{2}\right]}{\max(\eta,\gamma)^{2}}+T\delta$.
@@
By the law of total expectation, the fact that $|\mathcal{T}_{E}|\leq T$ and Lemma 5.4 that $\mathbb{P}\{\mathcal{E}^{c}\}\leq\delta$,
$\mathbb{E}\left[|\mathcal{T}_{E}|\right]=\mathbb{E}\left[|\mathcal{T}_{E}|\,\middle|\,\mathcal{E}\right]\mathbb{P}\{\mathcal{E}\}+\mathbb{E}\left[|\mathcal{T}_{E}\,\middle|\,\mathcal{E}^{c}|\right]\mathbb{P}\{\mathcal{E}^{c}\}\leq\mathbb{E}\left[|\mathcal{T}_{E}|\,\middle|\,\mathcal{E}\right]+T\delta.$ (23)
@@
It remains to upper bound $\mathbb{E}\left[|\mathcal{T}_{E}|\,\middle|\,\mathcal{E}\right]$. We condition on $\mathcal{E}$, which implies that $\boldsymbol{\theta}^{\star}\in\mathcal{C}_{t}$ for any $t\geq 0$. Focusing on posts that are admitted in the label-driven queue $t\in\mathcal{T}_{E}$ and letting $k=k(j(t))$ be the type of the post arrived in period $t$, it holds that $E_{k}(t)=1$. This also implies that $\underline{h}_{k}(t)<-\gamma$ and $\gamma<\bar{h}_{k}(t)$ by the label-driven admission rule (Line 3 of Algorithm 3). In addition, since $\boldsymbol{\theta}^{\star}\in\mathcal{C}_{t}$, it holds that $\underline{h}_{k}(t)\leq h_{k}(t)\leq\bar{h}_{k}(t)$. By the margin assumption $|h_{k}(t)|\geq\eta$, we must have $\bar{h}_{k}(t)-\underline{h}_{k}(t)\geq\eta+\gamma\geq\max(\eta,\gamma)$. To see this, if $h_{k}(t)\geq 0$, then $\bar{h}_{k}(t)\geq\eta$ and $\bar{h}_{k}(t)-\underline{h}_{k}(t)\geq\eta-(-\gamma)=\eta+\gamma$; same analysis holds for $h_{k}(t)<0$. In addition, for any $\boldsymbol{\theta}\in\mathcal{C}_{t-1}$,
$\left|\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}-\boldsymbol{\phi}_{k}^{\intercal}\hat{\boldsymbol{\theta}}_{t-1}\right|=\left|\boldsymbol{\phi}_{k}^{\intercal}(\boldsymbol{\theta}-\hat{\boldsymbol{\theta}}_{t-1})\right|\leq\|\boldsymbol{\theta}-\hat{\boldsymbol{\theta}}_{t-1}\|_{\bar{\boldsymbol{V}}_{t-1}}\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}\leq B_{\delta}(t-1)\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}$
@@
where the first inequality is by Cauchy Inequality and the second inequality is by the definition of $\mathcal{C}_{t-1}$ in (14). Combining this result with the definition of confidence intervals in (15), we have
$\max(\eta,\gamma)\leq\bar{h}_{k}(t)-\underline{h}_{k}(t)\leq\max_{\boldsymbol{\theta}\in\mathcal{C}_{t-1}}\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}-\min_{\boldsymbol{\theta}\in\mathcal{C}_{t-1}}\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}\leq 2B_{\delta}(t-1)\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}},$
@@
so $4B_{\delta}^{2}(T)\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}^{2}\geq\max(\eta,\gamma)^{2}$ and $4B_{\delta}^{2}(T)\sum_{t\in|\mathcal{T}_{E}|}\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}^{2}\geq|\mathcal{T}_{E}|\max(\eta,\gamma)^{2}$ conditioned on $\mathcal{E}$. This shows that $\mathbb{E}\left[|\mathcal{T}_{E}|\,\middle|\,\mathcal{E}\right]\leq\frac{4B_{\delta}^{2}(T)\sum_{t\in|\mathcal{T}_{E}|}\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}^{2}}{\max(\eta,\gamma)^{2}}$, which finishes the proof by (23). ∎
@@
where the first inequality is by Fact 4 (with $\kappa\geq\max(1,U^{2})$); the second inequality is by Fact 5; the last inequality is by $\kappa\geq U^{2},M\leq T$. Applying Lemma D.1 gives $\mathbb{E}\left[|\mathcal{T}_{E}|\right]\leq\frac{8B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}}+T\delta\leq\frac{9B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}}$ where the last inequality is because $\max(\eta,\gamma)\leq 1$, $T\delta\leq 1\leq B_{\delta}^{2}(T)d\ln(1+T/d)$ as $B_{\delta}^{2}(T)\geq\kappa U^{2}\geq 1$ (we assume $U\geq 1$) and $d\ln(1+T/d)\geq d\ln(1+3/d)\geq 1$ by $T\geq 3$ and Fact 9. ∎
@@
We fix a period $t$ and a type $k$ such that a type-$k$ post arrives in period $t$ and it is not admitted into the label-driven queue, i.e., $\Lambda_{k}(t)=1$ and $E_{k}(t)=0$; otherwise $Z_{k}(t)=0$. Since $\mathcal{E}$ holds, we have $\boldsymbol{\theta}^{\star}\in\mathcal{C}_{t-1}$ and thus $h_{k}\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]$ for any period $t$. We consider the three possible cases when the algorithm makes the classification decision $Y_{k}(t)$:
@@
  • $\underline{h}_{k}(t)\geq-\gamma$: we first suppose the empirical estimate of expected cost $\boldsymbol{\phi}_{k}^{\intercal}\hat{\boldsymbol{\theta}}(t)>0$. In this case we classify the post by $Y_{k}(t)=0$ (Line 3 in Algorithm 3). Then, if $h_{k}>0$, we have $Z_{k}(t)=0$ since $r_{k}^{R}\leq r_{k}^{O}$ by (5). If $h_{k}\leq 0$, we have $-\gamma\leq\underline{h}_{k}(t)\leq h_{k}\leq 0$ and $Z_{k}(t)\leq 3\gamma v_{k}\ell_{k}$ as in (20).
@@
    We next consider the case where the empirical estimate of expected cost $\boldsymbol{\phi}_{k}^{\intercal}\hat{\boldsymbol{\theta}}(t)\leq 0$. We argue that we must have $\bar{h}_{k}(t)\leq\gamma$. In this case, since $h_{k}\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]\subseteq[-\gamma,\gamma]$, we also have $Z_{k}(t)\leq 3\gamma v_{k}\ell_{k}$ by the same argument in (20). It remains to show $\bar{h}_{k}(t)\leq\gamma$. Suppose not, then there exists $\boldsymbol{\theta}\in\mathcal{C}_{t-1}$ such that $\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}>\gamma$. In addition, we know $\boldsymbol{\theta}^{\prime}=2\hat{\boldsymbol{\theta}}(t)-\boldsymbol{\theta}$ is also in $\mathcal{C}_{t-1}$ since $\boldsymbol{\theta}^{\prime}-\hat{\boldsymbol{\theta}}(t)=-(\boldsymbol{\theta}-\hat{\boldsymbol{\theta}}(t))$ and thus the matrix norms $\|\boldsymbol{\theta}^{\prime}-\hat{\boldsymbol{\theta}}(t)\|_{\bar{V}_{t-1}^{-1}},\|\boldsymbol{\theta}-\hat{\boldsymbol{\theta}}(t)\|_{\bar{V}_{t-1}^{-1}}$ are the same. However, we have $\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{\prime}=\boldsymbol{\phi}_{k}^{\intercal}(2\hat{\boldsymbol{\theta}}(t)-\boldsymbol{\theta})<0-\gamma=-\gamma$, which contradicts to the assumption that $\underline{h}_{k}(t)=\min_{\boldsymbol{\theta}_{1}\in\mathcal{C}_{t-1}}\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}_{1}\geq-\gamma$. Therefore, we must have $\bar{h}_{k}(t)\leq\gamma$. Summarizing the two cases shows that if $\underline{h}_{k}(t)\geq-\gamma$ then $Z_{k}(t)\leq 3\gamma v_{k}\ell_{k}.$
  • $\bar{h}_{k}(t)\leq\gamma$: we also have $Z_{k}(t)\leq 3\gamma v_{k}\ell_{k}$ by a symmetric argument of the above case;
@@
Combining the above cases, $Z_{k}(t)\leq(3\gamma v_{\max}\ell_{\max}+r_{\max}\ell_{\max}Q^{\mathcal{\textsc{LD}}}(t))\Lambda_{k}(t)$ conditioned on $\mathcal{E}$. ∎
@@
The proof is close to that of Lemmas 3.3 and 4.10, but relies on the analysis of a different Lyapunov function.
@@
$\leq 2\beta r_{\max}\ell_{\max}\mathbb{E}\left[\sum_{t=1}^{T}Q^{\mathcal{\textsc{LD}}}(t)\right]\leq\frac{18\beta r_{\max}\ell_{\max}B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}},$

### v3

@@ D.1 Comparison with Literature of Linear Contextual Bandits
Contrasting to linear contextual bandit, e.g. [APS11], and linear contextual bandit with delays, such as [VCP17, BXZ23], there are four main distinctions.
The first distinction is that our work needs to estimate the cost of keeping a post ($r_{k}^{O}=\mathbb{E}_{k}\left[c_{j}^{+}\right]$) and the cost of removing a post ($r_{k}^{R}=\mathbb{E}_{k}\left[-c_{j}^{-}\right]$) while papers in multi-armed bandits only need to estimate the mean reward of an arm. The reason is that we need to identify the type with high error of classification by AI — in multi-armed bandits, one only cares about types with high average cost (or reward). Indeed, since $r_{k}^{O}-r_{k}^{R}=h_{k}$, our setting requires to learn more information of underlying cost distributions. This requirement adds challenge to our concentration bound analysis since it is apriori unclear if $c_{j}^{+}$ (or $c_{j}^{-}$) is still sub-Gaussian when the cost $c_{j}$ is sub-Gaussian. Fortunately, our result in Lemma C.3 affirmatively answers this question and shows that a Lipschitz function of a sub-Gaussian random variable is still sub-Gaussian but with twice variance proxy.
The second distinction is that those works use a different estimator $\hat{\boldsymbol{\theta}}$ by including all arrived data pointss into the dataset (unobserved labels are set as 00). In contract, our confidence set is constructed based on all reviewed posts in (14), which is a subset of all arrived posts endogenously determined by our algorithm.
The third distinction, as a result of the second, is that conditioned on the set of arrived posts, the matrix $\bar{V}_{t}$ in (14) is a fixed matrix in their settings, but is a random matrix in our setting because it is constructed from the set of reviewed posts. We thus must deal with intricacy of the randomness in the regression.
The fourth distinction lies in the analysis with queueing-delayed feedback. The analysis of prior work crucially relies on an independence assumption, i.e., the event of observing feedback for a post within a particular delay period is independent from other posts, which enables a concentration bound on the number of observed feedback. In our setting, this is no longer the case as the delay in one post implies that other posts wait in the queue and thus delays could correlate with each other. We thus resort to the more intuitive estimator and analyze it via properties of our queueing systems.
@@ D.2 Regret guarantee of COLBACID (Theorem 4)
The proof uses [APS11, Theorem 2], which we restate in Fact 3. Recall that $\mathcal{C}^{O}_{t}$ is the set of vector $\boldsymbol{\theta}^{O}$ covered in $\mathcal{C}_{t}$ and $\mathcal{C}^{R}_{t}$ is the set of vector $\boldsymbol{\theta}^{R}$ covered in $\mathcal{C}_{t}$ as defined in (15). The proof proceeds by showing that $\boldsymbol{\theta}^{\star,O}\in\mathcal{C}^{O}_{t}$ and $\boldsymbol{\theta}^{\star,R}\in\mathcal{C}^{R}_{t}$ with probability $1-2\delta$, which implies $\boldsymbol{\Theta}^{\star}\in\mathcal{C}_{t}$ by the definition of $\mathcal{C}_{t}$ in (15).
We first show that $\boldsymbol{\theta}^{\star,O}\in\mathcal{C}^{O}_{t}$ with probability $1-\delta$. Recall that the data-point for $\hat{\boldsymbol{\theta}}^{O}$ in period $t$ is given by $Z_{t}=c^{+}_{\mathcal{S}(t)}$. Define $\xi_{\mathcal{S}(t)}=c^{+}_{\mathcal{S}(t)}-\mathbb{E}_{k(\mathcal{S}(t))}\left[c^{+}_{\mathcal{S}(t)}\right]=c^{+}_{\mathcal{S}(t)}-\boldsymbol{\phi}_{k(\mathcal{S}(t))}^{\intercal}\boldsymbol{\theta}^{\star,O}$. To apply Fact 3, it suffices to show that the sequence $\{\xi_{\mathcal{S}(t)}\}$ is conditionally sub-Gaussian with variance proxy $2\sigma^{2}_{\max}$ for the filtration $\mathcal{F}_{t-1}=\sigma(\boldsymbol{X}_{1},\ldots,\boldsymbol{X}_{t},\xi_{\mathcal{S}(1)},\ldots,\xi_{\mathcal{S}(t-1)})$, i.e., $\mathbb{E}\left[e^{u\xi_{\mathcal{S}(t)}}\,\middle|\,\mathcal{F}_{t-1}\right]\leq\exp(u^{2}(2\sigma_{\max}^{2})/2)$ for any $u\in\mathbb{R}$. Given that we assume post costs are independent samples, it holds that $\mathbb{E}\left[e^{u\xi_{\mathcal{S}(t)}}\,\middle|\,\mathcal{F}_{t-1}\right]=\mathbb{E}\left[e^{u\xi_{\mathcal{S}(\tau)}}\right]$. Since we assume $c_{\mathcal{S}(t)}$ is sub-Gaussian with variance proxy $\sigma_{\max}^{2}$ conditioned on the type of the post $k(\mathcal{S}(t))=k$ and $\max(x,0)$ is a $1-$Lipschitz function, Lemma C.3 shows that $c_{\mathcal{S}(t)}^{+}$ is sub-Gaussian with variance proxy $2\sigma_{\max}^{2}$. As a result, $\xi_{\mathcal{S}(t)}$ is also sub-Gaussian with variance proxy $2\sigma_{\max}^{2}$ and $\mathbb{E}\left[e^{u\xi_{\mathcal{S}(t)}}\right]\leq\exp(u^{2}(2\sigma_{\max}^{2})/2)$, showing that thus $\{\xi_{\mathcal{S}(t)}\}$ is conditionally sub-Gaussian with variance proxy $2\sigma_{\max}^{2}$. Hence the conditions of Fact 3 are met. Applying Fact 3 and the definition of $\mathcal{C}_{t}$ then gives $\boldsymbol{\theta}^{\star,O}\in\mathcal{C}_{t}^{O}$ with probability at least $1-\delta$.
By symmetry, the above argument applies to $\boldsymbol{\theta}^{\star,R}$ and thus $\boldsymbol{\theta}^{\star,R}\in\mathcal{C}_{t}^{R}$ with probability at least $1-\delta$. Using union bound gives that $\boldsymbol{\Theta}^{\star}\in\mathcal{C}_{t}$ with probability at least $1-2\delta$. ∎
@@
For $\delta\in(0,1)$, it holds that $\mathbb{E}\left[|\mathcal{T}_{E}|\right]\leq\frac{16B_{\delta}^{2}(T)\mathbb{E}\left[\sum_{t\in\mathcal{T}_{E}}\|\boldsymbol{\phi}_{k(j(t))}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}^{2}\right]}{\max(\eta,\gamma)^{2}}+2T\delta$.
@@
By the law of total expectation, the fact that $|\mathcal{T}_{E}|\leq T$ and Lemma 5.4 that $\mathbb{P}\{\mathcal{E}^{c}\}\leq 2\delta$,
$\mathbb{E}\left[|\mathcal{T}_{E}|\right]=\mathbb{E}\left[|\mathcal{T}_{E}|\,\middle|\,\mathcal{E}\right]\mathbb{P}\{\mathcal{E}\}+\mathbb{E}\left[|\mathcal{T}_{E}\,\middle|\,\mathcal{E}^{c}|\right]\mathbb{P}\{\mathcal{E}^{c}\}\leq\mathbb{E}\left[|\mathcal{T}_{E}|\,\middle|\,\mathcal{E}\right]+2T\delta.$ (24)
@@
It remains to upper bound $\mathbb{E}\left[|\mathcal{T}_{E}|\,\middle|\,\mathcal{E}\right]$. We condition on $\mathcal{E}$, which implies that $\boldsymbol{\Theta}^{\star}\in\mathcal{C}_{t}$ for any $t\geq 0$. Focusing on posts that are admitted in the label-driven queue $t\in\mathcal{T}_{E}$ and letting $k=k(j(t))$ be the type of the post arrived in period $t$, it holds that $E_{k}(t)=1$. This also implies that $\underline{h}_{k}(t)<-\gamma$ and $\gamma<\bar{h}_{k}(t)$ by the label-driven admission rule (Line 3 of Algorithm 3). In addition, since $\boldsymbol{\Theta}^{\star}\in\mathcal{C}_{t}$, it holds that $\underline{h}_{k}(t)\leq h_{k}\leq\bar{h}_{k}(t)$. By the margin assumption $|h_{k}(t)|\geq\eta$, we must have $\bar{h}_{k}(t)-\underline{h}_{k}(t)\geq\eta+\gamma\geq\max(\eta,\gamma)$. To see this, if $h_{k}\geq 0$, then $\bar{h}_{k}(t)\geq\eta$ and $\bar{h}_{k}(t)-\underline{h}_{k}(t)\geq\eta-(-\gamma)=\eta+\gamma$; same analysis holds for $h_{k}<0$. In addition, recall $\hat{\boldsymbol{\theta}}^{h}(t-1)=\hat{\boldsymbol{\theta}}^{O}(t-1)-\hat{\boldsymbol{\theta}}^{R}(t-1)$. For any $\boldsymbol{\theta}^{h}\in\mathcal{C}^{h}_{t-1}$, there exists $\boldsymbol{\theta}^{O}\in\mathcal{C}^{O}_{t-1}$ and $\boldsymbol{\theta}^{R}\in\mathcal{C}^{R}_{t-1}$ such that $\boldsymbol{\theta}^{h}=\boldsymbol{\theta}^{O}-\boldsymbol{\theta}^{R}$. Furthermore,
$\left|\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{h}-\boldsymbol{\phi}_{k}^{\intercal}\hat{\boldsymbol{\theta}}^{h}_{t-1}\right|=\left|\boldsymbol{\phi}_{k}^{\intercal}(\boldsymbol{\theta}^{h}-\hat{\boldsymbol{\theta}}^{h}_{t-1})\right|$ $\leq\|\boldsymbol{\theta}^{h}-\hat{\boldsymbol{\theta}}^{h}_{t-1}\|_{\bar{\boldsymbol{V}}_{t-1}}\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}$
$=\|\boldsymbol{\theta}^{O}-\boldsymbol{\theta}^{R}-(\hat{\boldsymbol{\theta}}^{O}(t-1)-\hat{\boldsymbol{\theta}}^{R}(t-1))\|_{\bar{\boldsymbol{V}}_{t-1}}\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}$
$\leq\left(\|\boldsymbol{\theta}^{O}-\hat{\boldsymbol{\theta}}^{O}(t-1)\|_{\bar{\boldsymbol{V}}_{t-1}}+\|\boldsymbol{\theta}^{R}-\hat{\boldsymbol{\theta}}^{R}(t-1)\|_{\bar{\boldsymbol{V}}_{t-1}}\right)\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}$
$\leq 2B_{\delta}(t-1)\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}$
@@
where the first inequality is by Cauchy Inequality; the second inequality is by triangle inequality; and the last inequality is by the definition of $\mathcal{C}^{O}_{t-1},\mathcal{C}^{R}_{t-1}$ in (15). Combining this result with the definition of confidence intervals in (16), we have
$\max(\eta,\gamma)\leq\bar{h}_{k}(t)-\underline{h}_{k}(t)\leq\max_{\boldsymbol{\theta}^{h}\in\mathcal{C}^{h}_{t-1}}\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}-\min_{\boldsymbol{\theta}^{h}\in\mathcal{C}^{h}_{t-1}}\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}\leq 4B_{\delta}(t-1)\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}},$
@@
so $16B_{\delta}^{2}(T)\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}^{2}\geq\max(\eta,\gamma)^{2}$ and $16B_{\delta}^{2}(T)\sum_{t\in|\mathcal{T}_{E}|}\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}^{2}\geq|\mathcal{T}_{E}|\max(\eta,\gamma)^{2}$ conditioned on $\mathcal{E}$. This shows that $\mathbb{E}\left[|\mathcal{T}_{E}|\,\middle|\,\mathcal{E}\right]\leq\frac{16B_{\delta}^{2}(T)\sum_{t\in|\mathcal{T}_{E}|}\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}^{2}}{\max(\eta,\gamma)^{2}}$, which finishes the proof by (24). ∎
@@
where the first inequality is by Fact 4 (with $\kappa\geq\max(1,U^{2})$); the second inequality is by Fact 5; the last inequality is by $\kappa\geq U^{2},M\leq T$. Applying Lemma D.1 gives $\mathbb{E}\left[|\mathcal{T}_{E}|\right]\leq\frac{32B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}}+2T\delta\leq\frac{34B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}}$ where the last inequality is because $\max(\eta,\gamma)\leq 1$, $T\delta\leq 1\leq B_{\delta}^{2}(T)d\ln(1+T/d)$ as $B_{\delta}^{2}(T)\geq\kappa U^{2}\geq 1$ (we assume $U\geq 1$) and $d\ln(1+T/d)\geq d\ln(1+3/d)\geq 1$ by $T\geq 3$ and Fact 9. ∎
@@
We fix a period $t$ and a type $k$ such that a type-$k$ post arrives in period $t$ and it is not admitted into the label-driven queue, i.e., $\Lambda_{k}(t)=1$ and $E_{k}(t)=0$; otherwise $Z_{k}(t)=0$. Since $\mathcal{E}$ holds, we have $\boldsymbol{\theta}^{\star}_{h}\in\mathcal{C}^{h}_{t-1}$ and thus $h_{k}\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]$ for any period $t$. We consider the three possible cases when the algorithm makes the classification decision $Y_{k}(t)$:
@@
  • $\underline{h}_{k}(t)\geq-\gamma$: we first suppose the empirical estimate of expected cost $\boldsymbol{\phi}_{k}^{\intercal}\hat{\boldsymbol{\theta}}^{h}(t-1)>0$ where $\hat{\boldsymbol{\theta}}^{h}(t-1)=\hat{\boldsymbol{\theta}}^{O}(t-1)-\hat{\boldsymbol{\theta}}^{R}(t-1)$. In this case we classify the post by $Y_{k}(t)=0$ (Line 3 in Algorithm 3). Then, if $h_{k}>0$, we have $Z_{k}(t)=0$ since $r_{k}^{R}=r_{k}^{O}-h_{k}\leq r_{k}^{O}$. If $h_{k}\leq 0$, we have $-\gamma\leq\underline{h}_{k}(t)\leq h_{k}\leq 0$ and $Z_{k}(t)\leq\gamma\ell_{k}$ as in (22).
@@
    We next consider the case where the empirical estimate of expected cost $\boldsymbol{\phi}_{k}^{\intercal}\hat{\boldsymbol{\theta}}^{h}(t-1)\leq 0$. We argue that we must have $\bar{h}_{k}(t)\leq\gamma$. In this case, since $h_{k}\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]\subseteq[-\gamma,\gamma]$, we also have $Z_{k}(t)\leq\gamma\ell_{k}$ by the same argument in (22). It remains to show $\bar{h}_{k}(t)\leq\gamma$. Suppose not, then there exists $\boldsymbol{\theta}^{h}\in\mathcal{C}^{h}_{t-1}$ such that $\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{h}>\gamma$. By definition of $\mathcal{C}^{h}_{t-1}$, there exists $\boldsymbol{\theta}^{O}\in\mathcal{C}^{O}_{t-1}$ and $\boldsymbol{\theta}^{R}\in\mathcal{C}^{R}_{t-1}$ such that $\boldsymbol{\theta}^{h}=\boldsymbol{\theta}^{O}-\boldsymbol{\theta}^{R}$. Define $\boldsymbol{\theta}^{{}^{\prime},O}\coloneqq 2\hat{\boldsymbol{\theta}}^{O}(t-1)-\boldsymbol{\theta}^{O},\boldsymbol{\theta}^{{}^{\prime},R}\coloneqq 2\hat{\boldsymbol{\theta}}^{R}(t-1)-\boldsymbol{\theta}^{R}$ and $\boldsymbol{\theta}^{{}^{\prime},h}\coloneqq\boldsymbol{\theta}^{{}^{\prime},O}-\boldsymbol{\theta}^{{}^{\prime},R}.$ We know $\boldsymbol{\theta}^{{}^{\prime},O}$ is in $\mathcal{C}^{O}_{t-1}$ since $\boldsymbol{\theta}^{{}^{\prime},O}-\hat{\boldsymbol{\theta}}^{O}(t-1)=-(\boldsymbol{\theta}^{{}^{\prime},O}-\hat{\boldsymbol{\theta}}(t-1))$ and thus the matrix norms $\|\boldsymbol{\theta}^{{}^{\prime},O}-\hat{\boldsymbol{\theta}}^{O}(t-1)\|_{\bar{V}_{t-1}^{-1}},\|\boldsymbol{\theta}^{O}-\hat{\boldsymbol{\theta}}^{O}(t-1)\|_{\bar{V}_{t-1}^{-1}}$ are the same. Similarly we have $\boldsymbol{\theta}^{{}^{\prime},R}\in\mathcal{C}^{R}_{t-1}$ and thus $\boldsymbol{\theta}^{{}^{\prime},h}\in\mathcal{C}^{h}_{t-1}.$ However, we have $\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{{}^{\prime},h}=\boldsymbol{\phi}_{k}^{\intercal}(2\hat{\boldsymbol{\theta}}^{h}(t-1)-\boldsymbol{\theta}^{h})<0-\gamma=-\gamma$, which contradicts to the assumption that $\underline{h}_{k}(t)=\min_{\boldsymbol{\theta}\in\mathcal{C}^{h}_{t-1}}\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}\geq-\gamma$. Therefore, we must have $\bar{h}_{k}(t)\leq\gamma$. Summarizing the two cases shows that if $\underline{h}_{k}(t)\geq-\gamma$ then $Z_{k}(t)\leq\gamma\ell_{k}.$
  • $\bar{h}_{k}(t)\leq\gamma$: we also have $Z_{k}(t)\leq\gamma\ell_{k}$ by a symmetric argument of the above case;
@@
Combining the above cases, $Z_{k}(t)\leq(\gamma\ell_{\max}+r_{\max}\ell_{\max}Q^{\mathcal{\textsc{LD}}}(t))\Lambda_{k}(t)$ conditioned on $\mathcal{E}$. ∎
@@
The proof is close to that of Lemmas 3.3 and 4.9, but relies on the analysis of a different Lyapunov function.
@@
$\leq 2\beta r_{\max}\ell_{\max}\mathbb{E}\left[\sum_{t=1}^{T}Q^{\mathcal{\textsc{LD}}}(t)\right]\leq\frac{68\beta r_{\max}\ell_{\max}B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}},$